The fake monster formal group. 27 May, 31 Dec 1998

Duke Math J. Vol 100 No. 1 (1999), 139-165.

Richard E. Borcherds, ^{∗}

D.P.M.M.S., 16 Mill Lane, Cambridge, CB2 1SB, England.

e-mail: reb@dpmms.cam.ac.uk

home page: www.dpmms.cam.ac.uk/˜ reb

Contents.

1. Introduction.

Notation and terminology.

2. Some theorems about smooth Hopf algebras.

3. Liftings of Lie algebra elements.

4. The smooth fake monster Hopf algebra.

5. A smooth Hopf algebra for the Virasoro algebra.

6. The no-ghost theorem over Z.

7. An application to modular moonshine.

8. Open problems.

1. Introduction.

The main result of this paper is the construction of “good” integral forms for the universal enveloping algebras of the fake monster Lie algebra and the Virasoro algebra. As an application we construct formal group laws over the integers for these Lie algebras. We also prove a form of the no-ghost theorem over the integers, and use this to verify an assumption used in the proof of the modular moonshine conjectures.

Over the integers the universal enveloping algebra of a Lie algebra is not very well behaved, and it is necessary to use a better integral form of the universal enveloping algebra over the rational numbers. The correct notion of “good” integral form was found by Kostant [K]. He found that the good integral forms are the ones with a structural base (as defined in 2.3), and showed that the universal enveloping algebras of finite dimensional semisimple Lie algebras have a structural base. The existence of a structural base implies that the dual algebra of the underlying coalgebra is a ring of formal power series. As this ring can be thought of as a sort of “coordinate ring” of some sort of formal group, this condition can be thought of as saying that the formal group is smooth and connected, and that the formal group comes from a “formal group law”.

The main point of this paper is to find such integral forms for the universal enveloping algebras of certain infinite dimensional Lie algebras. We recall that the universal enveloping algebra of a Lie algebra has a natural structure of a cocommutative Hopf algebra. So we need some theorems to tell us when a Hopf algebra has a structural base. In section 2 we prove that a Hopf algebra H has a structural base provided that there are “sufficiently many” group-like elements 1 + a_{1}x + a_{2}x^{2} + ⋯ ∈ H[[x]], and provided a few other minor conditions are satisfied. Such a group-like element should be thought of as an infinitesimal curve in the formal group, so roughly speaking this condition means that a set of generators

∗ Supported by a Royal Society professorship and by an NSF grant. of the Lie algebra of the formal group can be lifted to formal curves in the formal group. As an example of the theorems in section 2, we find that the Lie bracket [a_{1}, b_{1}] of two primitive liftable elements is also liftable. At first sight this looks as if it should be easy to prove, as all we need to do is write down an explicit lifting, with coefficients that are some universal (non-commutative) polynomials in the coefficients of liftings of a_{1} and b_{1}. However this seems to be very hard to do explicitly (partly because such a lifting is far from unique, which paradoxically makes it harder to find one). Instead, we use a far more roundabout argument, where we first need to prove a theorem (theorem 2.12) saying that all primitive elements of certain Hopf algebras are liftable.

We also give an incidental application of these theorems to certain Hopf algebras F_{n} considered by Dieudonné. He showed that these Hopf algebras have a structural basis when reduced mod p for any prime p; we show that the Hopf algebras F_{n} already have a structural basis over the integers.

In section 2 we have reduced the problem of finding a structural basis of a Hopf algebra H to the problem of finding enough group-like elements in H[[x]]. For the universal enveloping algebras of finite dimensional simple Lie algebras, or more generally for Kac-Moody algebras, this is easy to do, because we just take the formal 1-parameter subgroup

∑_{n≥0} x^{n}Ad(e)^{n}/n!

for (locally nilpotent) elements e of the Lie algebra corresponding to real roots, and define the integral form of the Hopf algebra to be the one generated by the coefficients of these liftings. This gives the usual Kostant integral form for the universal enveloping algebra of finite dimensional semisimple Lie algebras. For the Lie algebras in this paper this does not work because there are not enough locally nilpotent elements. (We can try to use the elements e^{n}/n! for non nilpotent elements, but there seems no obvious way in which this gives a good integral form.) Fortunately we do not need to lift the generators of the Lie algebra to formal one parameter groups, and it is sufficient to lift them to formal curves, which is easier to do.

In section 3 we construct liftings of some elements of the vertex algebra of a lattice to formal curves. More precisely, we show that we can lift any element of a certain Lie algebra provided that it lies in a root space of a root of norm 2 or 0. We do this by explicitly writing down a group-like lifting, and checking by brute force that its coefficients are integral. (Unfortunately there seems no obvious way to extend this brute force approach to root spaces of negative norm roots.)

In section 4 we use the liftings of section 3 to construct a smooth integral form of the universal enveloping algebra of the fake monster Lie algebra, which is constructed from the lattice II_{25,1}. The main point about this Lie algebra is that it is generated by the root spaces of norm 2 and norm 0 roots, so in section 3 we have constructed enough liftings to apply the theorems in section 2. (This is a very special property of II_{25,1}; it is the only known indefinite lattice whose Lie algebra is generated by the root spaces of norm 2 and norm 0 roots. In other words the theory in sections 2 and 3 has been developed mainly for this one example!) We can summarize the main results about the fake monster Lie algebra proved in this paper as follows.

Theorem 4.1. There is a II_{25,1}-graded Hopf algebra U^{+}(m) over Z with the following properties.

1 U^{+}(m) has a structural basis over Z.

2 The primitive elements of U^{+}(m) are an integral form of the fake monster Lie algebra

m. 3 For every norm 2 vector of II_{25,1}, U^{+}(m) contains the usual (Kostant) integral form

of the universal enveloping algebra of the corresponding sl_{2}(Z).

In section 5 we construct a smooth integral form for the universal enveloping algebra of the Virasoro algebra, by applying the theorems in section 2 to a set of explicit liftings of the basis elements of the Virasoro algebra. In other words we construct a formal group law for the Virasoro algebra over the integers.

The no-ghost theorem in string theory states that a certain real vector space of states is positive definite (so it contains no negative norm vectors, which are sometimes called ghosts and which would prevent the space from being a Hilbert space). This real vector space has a natural integral form which can be made into a positive definite lattice using the inner product, and we can ask for an integral form of the no-ghost theorem, which should say something about the structure of this lattice. In section 6 we use the smooth formal group of the Virasoro algebra to prove an integral form of the no-ghost theorem, at least in the case of vertex algebras constructed from 26 dimensional lattices. More precisely we prove bounds on the determinant of this lattice, which quite often imply that the lattice is self dual.

The proof [B98] of Ryba’s modular moonshine conjectures for large primes used an unproved technical assumption about the monster Lie algebra. In section 7 we use the integral no-ghost theorem to prove this assumption, thus completing the proof of the modular moonshine conjectures for odd primes. (The proof in [B-R] for the prime 2 still relies on another unproved technical assumption.)

Notation and terminology.

α An element of a lattice.

β An element of a lattice.

γ A norm 0 element of a lattice.

Γ An indeterminate.

c An element generating the center of the Virasoro algebra. Der The derivations of a commutative ring. e^{α} An element of the group ring of a lattice L, usually regarded as an element of the

vertex algebra of L. ϵ_{i} The function from I to Z taking value 1 on i ∈ I and 0 elsewhere. F_{p} A finite field of order p.

F (λ) If λ = 1^{i_{1}}2^{i_{2}} ⋯ is a partition, then F (λ) = i_{1}!i_{2}! · · ·. γ A norm 0 vector of a lattice L. H A cocommutative Hopf algebra. I An index set, or a finite sequence of integers.

II_{m,n} An even self dual lattice of dimension m + n and signature m − n. J A finite sequence of integers. l(I) The length of the sequence I. l(λ) If λ = 1^{i_{1}}2^{i_{2}} ⋯ is a partition, then l(λ) = i_{1} + i_{2} + · · ·.

λ A partition. K A field, sometimes the quotient field of R. L A lattice, usually even and self dual. L_{i} A basis element for the Virasoro or Witt algebra. m An integral form of the fake monster Lie algebra (or the monster Lie algebra in section

7). N An integer.

P (λ) If λ = 1^{i_{1}}2^{i_{2}} ⋯ is a partition, then P (λ) is the integer 1^{i_{1}}2^{i_{2}} · · ·.

p A prime. p(n) The number of partitions of n.

Q The rational numbers. ρ A Weyl vector. R A commutative ring. S The antipode of a Hopf algebra.

Σ(I) The sum of the elements of I. U A universal enveloping algebra. U^{+} An integral form of a universal enveloping algebra with a structural basis. V A vertex algebra. V_{p} The shift (or Verschiebung) of a coalgebra over F_{p}. V ir The Virasoro algebra. W itt The Witt algebra.

x A formal variable. Z The integers. Z_{≥0} is the set of non-negative integers. Z_{p} The p-adic integers. Z_{α} An element of a structural basis. z A formal variable.

Bialgebra A module with compatible algebra and coalgebra structures. Group-like Δ(a) = a ⊗ a. Hopf algebra A bialgebra with antipode. Liftable See definition 2.2. Primitive Δ(a) = a ⊗ 1 + 1 ⊗ a. Smooth. A Hopf algebra or coalgebra with a structural basis. Structural basis ΔZ_{α} = ∑_{0≤β≤α} Z_{β} ⊗ Z_{α−β}

2. Some theorems about Hopf algebras.

In this section we will prove several results about Hopf algebras. The main result is theorem 2.15, which states that under mild conditions a bialgebra generated by the coefficients of group-like liftings has a structural basis. In later sections we will use this to construct a good integral form of the universal enveloping algebra of the fake monster Lie algebra.

We recall that a coalgebra over a commutative ring R is an R-module H with a coassociative coproduct and a counit, a bialgebra is a module with compatible algebra and coalgebra structures, and a Hopf algebra is a bialgebra with an antipode S.

Definition 2.1. An element a of a coalgebra is called group-like if Δ(a) = a ⊗ a, and primitive if Δ(a) = a ⊗ 1 + 1 ⊗ a.

Definition 2.2. We say that an element a_{1} of the coalgebra H is liftable to a group-like element, or liftable for short, if we can find elements a_{n} ∈ H (n ≥ 0) such that a_{0} = 1 and the element ∑_{n≥0} a_{n}z^{n} ∈ H[[z]] is group-like. In other words

Δ(a_{n}) = ∑_{0≤m≤n} a_{m} ⊗ a_{n−m}

for all n ≥ 0. We say that a_{1} is liftable to order N if we can find elements a_{n} for 0 ≤ n ≤ N (with a_{0} = 1) satisfying the relations above for 0 ≤ n ≤ N. We say that the lifting is graded if H is graded by some abelian group L and deg(a_{n}) = nα for some α ∈ L.

We write Z_{≥0} for the nonnegative integers, and Z_{≥0}^{(I)} for the functions from a set I to

Z_{≥0} that are zero on all but a finite number of elements of I. The element ϵ_{i} of Z_{≥0}^{(I)} is defined to be the function that is 1 on i ∈ I and 0 elsewhere. We define a partial order ≥ on Z_{≥0}^{(I)} in the obvious way, by saying α ≥ β if α(i) ≥ β(i) for all i ∈ I.

Definition 2.3. A structural basis for a coalgebra or Hopf algebra H over a ring R is a set of elements Z_{α}, α ∈ Z_{≥0}^{(I)} for some set I, such that the elements Z_{α} form a basis for the free R-module H and such that ∑_{α} Z_{α}x^{α} is group-like, in other words

Δ(Z_{α}) = ∑_{0≤β≤α} Z_{β} ⊗ Z_{α−β}.

Lemma 2.4. If a coalgebra H has a structural basis Z_{α} then the R-module of primitive elements of H has a basis consisting of the elements Z_{ϵ_{i}} for i ∈ I.

Proof. We have to show that every primitive element of H is a linear combination of the elements Z_{ϵ_{i}}. Give H the bialgebra structure such that Z_{α}Z_{β} = Z_{α+β} ∏_{i∈I} \binom{α(i)+β(i)}{β(i)} (as in [A, section 2.5.1] when R is a field). The dual algebra H^{∗} is a ring of formal power series and H acts on H^{∗} as differential operators. Suppose D is any primitive element of H. By subtracting multiples of Z_{ϵ_{i}} from D we can assume that D acts trivially on the elements x_{i} dual to Z_{ϵ_{i}}. But then the fact that D is a derivation implies that D acts trivially on any polynomial in the x_{i}’s. As there are no elements of H orthogonal to all polynomials in the x_{i}’s this proves that D must be 0. This proves lemma 2.4.

Lemma 2.5. The liftable elements of a bialgebra over R form an R-submodule.

Proof. If a(x) and b(x) are lifts of a_{1} and b_{1} then a(x)b(x) is a lift of a_{1} + b_{1}, and if r ∈ R then a(rx) is a lift of ra_{1}. This proves lemma 2.5.

Lemma 2.6. If H is a bialgebra with a structural basis then all primitive elements of H are liftable.

Proof. By lemma 2.4 the R-module of primitive elements has a base of elements Z_{ϵ_{i}}. Each of these basis elements can be lifted by ∑_{n≥0} Z_{nϵ_{i}}x^{n}. Lemma 2.6 now follows from lemma 2.5.

Lemma 2.7. Suppose H is a coalgebra with a structural basis over an integral domain R. Then any nonzero coideal J of H contains a nonzero primitive element of H.

Proof. We first assume that R is a field. A coalgebra over a field with a structural basis is irreducible and pointed (by [A, section 2.5]), so lemma 2.7 follows immediately from [A, corollary 2.4.14], which states that any nonzero coideal of a pointed irreducible coalgebra over a field contains a nonzero primitive element.

For the general case, let K be the quotient field of R. If the coideal J is nonzero then J ⊗_{R} K is a nonzero coideal of H ⊗_{R} K. As lemma 2.7 is is true over the field K this shows that J ⊗_{R} K contains a primitive element, and multiplying this by a suitable nonzero constant to cancel the denominators gives a nonzero primitive element of J. This proves lemma 2.7.

Lemma 2.8. Suppose that H is a torsion-free coalgebra over Z such that for every prime p every primitive element of H/pH is the image of a primitive element of H. Let N be any positive integer. Then every primitive element of H/NH is the image of a primitive element of H.

Proof. We prove this by induction on the number of prime factors of N, the case N = 1 being trivial. Suppose p is any prime factor of N, and suppose that the image of x ∈ H is primitive in H/NH, so that we have to show that x is congruent to a primitive element mod N. The element x maps to a primitive element in H/pH, so by assumption we have x = y+pz for some primitive element y and some z. But then Δ(pz) ≡ pz⊗1+1⊗pz mod N as x and y are both primitive mod N. Using the fact that H is torsion-free we can divide through by p to find that Δ(z) ≡ z ⊗ 1 + 1 ⊗ z mod N/p. By induction on the number of prime factors of N this shows that z = t+(N/p)u for some primitive element t. Substituting this back in to x = y + pz shows that x = y + pt + Nu, so x is congruent to a primitive element y + pt mod N. This proves lemma 2.8.

Lemma 2.9. If every primitive element of a bialgebra is liftable to all finite orders, then every primitive element is liftable (to infinite order).

Note that we do not claim that if one particular primitive element of a cocommutative Hopf algebra can be lifted to all finite orders then it can be lifted to infinite order. I do not know whether or not this is true in general.

Proof. Suppose that 1 + a_{1}x + a_{2}x^{2} + ⋯ + a_{n}x^{n} is an order n lifting of a_{1}. It is sufficient to show that we can extend this to an order n + 1 lifting, because by repeating this an infinite number of times we get a lifting of a_{1} of infinite order. We will prove by induction on k that if k ≤ n then we can find an order n + 1 lifting of a_{1} which agrees with a(x) to order k. This is true for k = 1 by the assumption that a_{1} has liftings of arbitrary finite order. Let b(x) = 1 + b_{1}x + ⋯ + b_{n+1}x^{n+1} ∈ H[x]/(x^{n+2}) be an order n + 1 lifting of a_{1} with b_{i} = a_{i} for i ≤ k. Then a(x)b(x)^{−1} = 1 + (a_{k+1} − b_{k+1})x^{k+1} + ⋯ is group-like, so a_{k+1} − b_{k+1} is primitive. (Note that b(x)^{−1} is a well defined order n + 1 lifting, as we can construct it as ∑_{m}(1 − b(x))^{m}.) Let c(x) be an order n + 1 lifting of a_{k+1} − b_{k+1} (which exists as we assumed all primitive elements have liftings of any finite order). Then b(x)c(x^{k+1}) is an order n + 1 lifting of a_{1}, which agrees with a(x) to order k + 1. This shows that any order n lifting of a_{1} extends to an order n + 1 lifting. This proves lemma 2.9.

Lemma 2.10. Suppose that H is a torsion-free bialgebra over Z and suppose that for every prime p the map from primitive elements of H to primitive elements of H/pH is onto. Then every primitive element of H is liftable.

Proof. We will show by induction on n that any primitive element a can be lifted to a group-like element in H[x]/(x^{n+1}); this is trivial for n = 0. We will show by induction on k that if 1 ≤ k ≤ n then we can find a lifting ∑_{0≤i≤n} c_{i}x^{i} of n!a such that n!^{i}|c_{i} for i ≤ k. This is true for k = 1, because we can use the order n lifting ∑_{0≤i≤n}(n!^{i}/i!)a^{i}x^{i}. So suppose that k ≥ 2 and that n!^{i}|c_{i} for i < k. Then

Δ(c_{k}) = ∑_{0≤i≤k} c_{i} ⊗ c_{k−i} ≡ c_{k} ⊗ 1 + 1 ⊗ c_{k} mod n!^{k}.

Therefore c_{k} is primitive in H/n!^{k}H. By lemma 2.8 and the assumption that for every prime p the map from primitive elements of H to primitive elements of H/pH is onto this implies that c_{k} = d_{1} + n!^{k}e for some primitive d_{1} and some e. By induction on n we can find a lifting d(x) = ∑_{0≤i<n} d_{i}x^{i} ∈ H[x]/(x^{n}) of d_{1}, and changing x to x^{k} we get a group-like element d(x^{k}) ∈ H[x]/(x^{n+1}). (Here we use the fact that k ≥ 2 and n ≥ 1, so that kn ≥ n + 1.) Multiplying ∑ c_{i}x^{i} by d(x)^{−1} mod x^{n+1} we get a group-like element in H[x]/(x^{n+1}) whose coefficient of x^{i} for i ≤ k is divisible by n!^{k}. This completes the proof by induction on k, and shows that we can assume that n!^{i}|c_{i} for all i ≤ n. But then using the fact that H is torsion free we see that ∑(c_{i}/n!^{i})x^{i} is a lifting of c_{1}/n! = a in H[x]/(x^{n+1}).

We have shown that any primitive element of H can be lifted to a group-like element of any finite order n. By lemma 2.9 this implies that every primitive element of H is liftable. This proves lemma 2.10.

We recall that for any cocommutative coalgebra or Hopf algebra H over F_{p} there is a homomorphism V_{p} from H to itself, called the shift (or Verschiebung), which is dual to the Frobenius map x ↦ x^{p} on the dual algebra H^{∗} (see for example [A section 5.4]). The shift V_{p} commutes with all homomorphisms of coalgebras. If H is a coalgebra over Z we also write V_{p} for the shift of H/pH.

Lemma 2.11. Suppose that the Z_{α}’s for α ∈ Z_{≥0}^{(I)} are elements of a coalgebra H over F_{p} satisfying the relations

Δ(Z_{α}) = ∑_{0≤β≤α} Z_{β} ⊗ Z_{α−β}

Then V_{p}(Z_{α}) = Z_{α/p}, where Z_{α/p} means 0 if p does not divided α.

Proof. If the Z_{α}’s are linearly independent and span H, or in other words if they form a structural basis, then the result follows from [A, theorem 2.5.9]. In general we can find a coalgebra with a structural basis together with a homomorphism mapping the elements of the structural basis to the Z_{α}’s. The result then follows because the shifts V_{p} commute with homomorphisms. This proves lemma 2.11.

Theorem 2.12. Suppose that H is a Z_{≥0}-graded cocommutative bialgebra over Z whose homogeneous pieces are finitely generated free abelian groups and such that the degree 0 piece is spanned by 1. Suppose that the shifts V_{p} : H/pH → H/pH are onto for all primes p. Then every primitive element of H is liftable.

Remark. In fact theorem 2.15 (whose proof uses theorem 2.12) implies the stronger result that H has a structural basis.

Proof. We first show that if K is any field then the K bialgebra C = H ⊗ K is irreducible, in other words it has only one simple subcoalgebra. By theorem 2.3.4 of [A], the simple subcoalgebras of a K-coalgebra C correspond to the maximal ideals of the dual algebra C^{∗} which are not dense, so it is sufficient to show that C^{∗} is a local ring. The elements of C^{∗} are infinite sums of the form c = c_{0} + c_{1} + ⋯ with c_{i} ∈ C_{i}^{∗} (where C_{i} is the degree i subspace of C). We show that the ideal M of all elements with c_{0} = 0 is the unique maximal ideal of C^{∗}, in other words that any element c with c_{0} ≠ 0 invertible. The space C_{0}^{∗} is isomorphic to the field K, so if c_{0} ≠ 0 then it is invertible, so we can recursively define an inverse d = d_{0} +d_{1} +⋯ to c by putting d_{0} = c_{0}^{−1}, d_{n} = −c_{0}^{−1}(d_{n−1}c_{1} +⋯+d_{0}c_{n}) for n > 0. This shows that C^{∗} is a local ring, so C is irreducible.

Next we check that H ⊗ F_{p} and H ⊗ Q have structural bases. By [A, theorem 2.4.24] any irreducible bialgebra over a field, in particular H ⊗F_{p}, is automatically a Hopf algebra. By [A,corollary 2.5.15] any cocommutative irreducible Hopf algebra over F_{p} such that V_{p} is onto, in particular H ⊗F_{p}, has a structural basis. Similarly theorem 2.5.3 of [A] states that any irreducible cocommutative Hopf algebra over a field of characteristic 0 is a universal enveloping algebra of its primitive elements. So H ⊗ Q is the universal enveloping algebra of its primitive elements and therefore also has a structural basis.

The next step is to show that any primitive element of H ⊗ F_{p} can be lifted to a primitive element of H. If m_{α} is the dimension of the space of primitive elements of degree α of a Z_{≥0}^{n}-graded Hopf algebra with a structural basis and n_{α} is the dimension of the space of all elements of degree α, then

∏_{α>0}(1 − x^{α})^{−m_{α}} = ∑_{α≥0} n_{α}x^{α}

by lemma 2.4. In particular the spaces of primitive elements of degrees α in H ⊗ F_{p} and H ⊗ Q have the same dimension, because it is obvious that the spaces of all elements in H ⊗ F_{p} and H ⊗ Q of degree α have the same dimension, and H ⊗ F_{p} and H ⊗ Q both have structural bases by the paragraph above. The space P_{α} of primitive elements of H of degree α also has the same rank as the space of primitive elements of H ⊗Q, and therefore the same rank as the space P_{α,p} of primitive elements of H ⊗ F_{p} of degree α. The map from P_{α}/pP_{α} to P_{α,p} is obviously injective and both spaces are finite dimensional vector spaces over F_{p} of the same dimension, so the map between them is also surjective, or in other words every primitive element of H ⊗F_{p} (of any degree α) can be lifted to a primitive element of H. By lemma 2.10 this implies every primitive element of H is liftable. This proves theorem 2.12.

Corollary 2.13. Let k and n be positive integers. If

a(x) = ∑_{i_{1},...,i_{k}≥0} a_{i_{1},...,i_{k}}x_{1}^{i_{1}} ⋯ x_{k}^{i_{k}}

is group-like, for a_{i_{1},...,i_{k}} in some bialgebra over a commutative ring R, and a_{0,...,0} = 1, a_{i_{1},...,i_{k}} = 0 for 0 < i_{1} + ⋯ + i_{k} < n, then a_{i_{1},...,i_{k}} is liftable for i_{1} + ⋯ + i_{k} = n.

Proof. It is sufficient to prove this for the universal example, where the bialgebra H is the free associative algebra over Z generated by a_{i_{1},...,i_{k}} for i_{1} + ⋯ + i_{k} ≥ n and the coalgebra structure is defined so that a(x) is group-like. Put a_{0} = 1, a_{i_{1},...,i_{k}} = 0 for 0 < i_{1} + ⋯ i_{k} < n, and grade H by letting a_{i_{1},...,i_{k}} have degree i_{1} + ⋯ + i_{k}. Then V_{p}(a_{i_{1}....,i_{k}}) = a_{i_{1}/p,...,i_{k}/p} by lemma 2.11, so V_{p} : H/pH ↦ H/pH is onto. By theorem 2.12 this implies that any primitive element of H, in particular a_{i_{1},...,i_{k}} for i_{1} + ⋯ + i_{k} < 2n, is liftable. This proves corollary 2.13.

Theorem 2.14. The set of liftable elements of a cocommutative bialgebra H over a commutative ring R is a Lie algebra over R.

Proof. We already know that the liftable primitive elements form a module over R by lemma 2.5. We have to show that the set of liftable elements is closed under the Lie bracket. Suppose a_{1} and b_{1} are two primitive elements with group-like lifts a(x) = ∑ a_{i}x^{i} and b(y) = ∑_{i} b_{i}y^{i}. It is sufficient to prove theorem 2.14 in the case of the universal example, when H is the free associative algebra over Z generated by the elements a_{i} and b_{i} for i > 0, with the coalgebra structure defined by the fact that a(x) and b(y) are group-like. We Z-grade H by giving a_{i} and b_{i} degree i ∈ Z. By lemma 2.11 the shifts V_{p} contain all the elements a_{i} and b_{i} in their images, so the shifts are onto as they are homomorphisms of F_{p}-algebras. Then by theorem 2.12 we see that all primitive elements of H, in particular [a_{1}, b_{1}], are liftable. This proves theorem 2.14.

Open problem: give a direct proof of theorem 2.14 by writing down an explicit lifting of [a_{1}, b_{1}] as a non-commutative polynomial in the a_{i}’s and the b_{i}’s. This seems surprisingly difficult, possibly because such a lifting is far from unique so it is hard to think of a canonical way to define it.

Theorem 2.15. Suppose that U is a torsion free bialgebra over a principal ideal domain R, and suppose that the Lie algebra of primitive elements of U is a free R-module. If U is generated as an algebra by all coefficients a_{i} of some set of group-like elements 1 + a_{1}x + a_{2}x^{2} + ⋯ ∈ U[[x]] then U is a Hopf algebra with a structural basis.

Proof. As R is a principal ideal domain, any submodule of a free module is free, and in particular the Lie algebra L of liftable primitive elements of U is a free R module. Suppose that I is a totally ordered set indexing a basis a_{1}^{i}, i ∈ I of L. (The superscript on a_{1}^{i} is just used as an index and has nothing to do with powers of a_{1}.) For each a_{1}^{i} chose a lifting a^{i}(x) = ∑_{n} a_{n}^{i}(x). Define H to be the R-module spanned by the elements of the form a_{n_{1}}^{i_{1}} ⋯ a_{n_{k}}^{i_{k}} with i_{1} < i_{2} ⋯ < i_{k}.

We need to check that the natural map from H ⊗_{R} H to U ⊗_{R} U is injective, so that H ⊗_{R} H can be regarded as a submodule of U ⊗_{R} U. As U is torsion-free, its submodule H is also torsion free and therefore flat (as R is a principal ideal domain). Hence H ⊗_{R} H ⊆ U ⊗_{R} U (as H ⊗_{R} H ⊆ U ⊗_{R} H ⊆ U ⊗_{R} U).

The spanning set of H of elements of the form a_{n_{1}}^{i_{1}} ⋯ a_{n_{k}}^{i_{k}} satisfies the relations of a structural basis, so in particular H is a sub coalgebra of U. This set of elements is linearly independent by lemma 2.7 because the primitive elements a_{1}^{i} in it are linearly independent, so it forms a structural basis for the coalgebra H. We will show that U has a structural basis by showing that it is equal to H. To do this we will first show that H is closed under multiplication, and then show that H contains a set of generators of the algebra U.

We define a filtration of H by letting H_{n} be the subspace of H spanned by all the monomials a_{n_{1}}^{i_{1}} ⋯ a_{n_{k}}^{i_{k}} with n_{1} +⋯ n_{k} ≤ n. We will show by induction on N that H_{m}H_{n} ⊆ H_{m+n} whenever m + n ≤ N; this is trivial for N = 1. To do this it is sufficient to show that for any r, s with r + s ≤ N we have a_{r}^{i}a_{s}^{j} ≡ a_{s}^{j}a_{r}^{i} mod H_{r+s−1} whenever i > j, and a_{r}^{i}a_{s}^{i} ≡ \binom{r+s}{r}a_{r+s}^{i} mod H_{r+s−1}.

We will show by induction on k that if 1 ≤ k ≤ N then there is a group-like element c(x, y) = ∑ c_{ij}x^{i}y^{j} such that c_{ij} ∈ H_{i+j} if i + j ≤ N and c_{ij} = b_{j}a_{i} for i + j ≤ k. This is obvious for k = 1, as we can take c_{ij} = a_{i}b_{j}. Suppose we have proved this for some value of k < N, so we wish to prove it for k + 1. Consider the group-like element b(y)a(x)c(x, y)^{−1}. The coefficients of x^{i}y^{j} of this are 0 for 0 < i + j ≤ k, so by corollary 2.13 the coefficients for i + j = k + 1 are all liftable. For each of them choose a lifting d_{i}(z), such that the coefficient of x^{j} of any of these liftings is in H_{j}. We can construct such liftings as products a^{i_{1}}(r_{1}x)a^{i_{2}}(r_{2}x) ⋯ if d_{i} = r_{1}a^{i_{1}} + r_{2}a^{i_{2}} + ⋯ with i_{1} < i_{2} < · · ·. Then the product

e(x, y) = ( ∏_{1≤i≤k} d_{i}(x^{i}y^{k+1−i}))c(x, y)

satisfies the conditions above for k + 1. In fact by construction we see that e_{ij} = b_{j}a_{i} for i + j ≤ k + 1. Also the coefficient of x^{i}y^{j} in e(x, y) is either already an element of H_{i+j} or it is a sum of products of elements in H_{i_{1}}, H_{i_{2}}, … with i_{1} + i_{2} + ⋯ < i + j. If i + j ≤ N then by the inductive assumption on N such product is already in H_{i+j−1} ⊆ H_{i+j}, so the coefficient of x^{i}y^{j} is in H_{i+j} provided i + j ≤ N. This proves the inductive hypothesis for all k ≤ N.

If we take k = N in the inductive hypothesis above, we now see that b_{j}a_{i} ∈ H_{i+j} for i + j = N. By induction on N we see that b_{j}a_{i} ∈ H_{i+j} for all i, j.

The proof that a_{i}a_{j} = \binom{i+j}{j}a_{i+j} mod H_{i+j−1} is similar, except that we use the grouplike element a(x + y) = ∑_{i,j} \binom{i+j}{i}a_{i+j}x^{i}y^{j} instead of b(x)a(y).

This shows that H is a subalgebra of U (and hence a Hopf algebra with a structural basis).

Finally we check that H contains the coefficients of any lifting a(x) of any primitive element a_{1} of U. We will show by induction on n that if n ≥ 1 then there is a lifting b(x) of a_{1} such that all coefficients of b(x) are in H and b_{i} = a_{i} for i ≤ n. This is clear for n = 1 as we just take any lifting of a_{1} with coefficients in H. Suppose that b_{i} = a_{i} for i ≤ n. Then b(x)^{−1}a(x) has all its coefficients of x^{i} vanishing for 0 < i ≤ n, so by corollary 2.13 (with k = 1) we see that the coefficient of x^{n+1} is liftable. Choose a lifting c(x) with coefficients in H. Then b(x)c(x^{n+1}) has all coefficients in H and the coefficients of x^{i} for i ≤ n + 1 are equal to a_{i}. This proves the inductive hypothesis for n + 1, and hence for all n. Therefore all coefficients of a(x) are in H.

We have shown that H is a bialgebra with a structural basis and that H contains a set of generators of the algebra U. Therefore H is equal to U, and U has a structural basis. This proves theorem 2.15.

Example 2.16. Suppose n is a nonnegative integer. The Z-Hopf algebra F_{n} is defined to the free associative algebra over Z on a set of elements Z_{α}, α ∈ Z_{≥0}^{n}, α ≠ 0, with the comultiplication defined by

Δ(Z_{α}) = ∑_{0≤β≤α} Z_{β} ⊗ Z_{α−β}

on the generators (where Z_{0} = 1). Then it follows immediately from theorem 2.15 that F_{n} has a structural basis. The fact that F_{n} ⊗ F_{p} has a structural basis was first proved by Dieudonné; more precisely it follows from the comments at the end of section 15 of [D], which show that F_{n}(F_{p}) has a structural basis, together with theorem 3 of [D], which shows that Dieudonné’s definition of F_{n}(F_{p}) is equivalent to F_{n} ⊗ F_{p}. There are proofs that F_{1} has a structural basis in [D72, proposition 2.2] and [Sh], but unfortunately the paper [Sh] seems to be unpublished, and the proof in [D72] has a gap (see [H p. 516]): in [D72, page 5, line 6], it is implicitly assumed that I_{k}/I_{k}^{2} is torsion free. (It follows from example 2.16 that this is indeed true, but it seems rather hard to prove.)

Example 2.17. Suppose that U is a torsion free bialgebra over a principal ideal domain R, and suppose that the Lie algebra of primitive elements of U is a free R-module. Then U has a subalgebra H with a structural basis containing all subalgebras with a structural base. The primitive elements of H are exactly the liftable primitive elements of U, and H is generated as an algebra by all coefficients of all liftings of primitive elements of U. This follows easily from theorem 2.15.

3. Liftings of Lie algebra elements.

In this section we construct liftings of certain elements of the Lie algebra of the vertex algebra of a double cover of an even lattice ([B86]). More precisely we construct liftings of elements in the root spaces of vectors of norms 2 or 0. This will be enough for applications to the fake monster Lie algebra, because its rational form is generated by the root spaces of roots of norms 2 or 0. For vectors in norm 2 root spaces we construct a formal one parameter group lifting any element in the root space. For norm 0 roots we cannot do this, but have to make do with lifting vectors to formal curves.

Lemma 3.1. Suppose a is an element of a vertex algebra, and define a^{k} for k ≥ 0 by a^{0} = 1, a^{k+1} = a_{−1}(a^{k}). Then

a_{n}^{k}(b) = ∑_{0≤j≤k} \binom{k}{j} ∑_{{i_{1},...,i_{j} <0 \atop i_{j+1},...,i_{k}≥0} ; i_{1}+···+i_{k}=n−k} a_{i_{1}} ⋯ a_{i_{k}}b

Proof. This follows by induction on n using the fact that

(a_{−1}a^{k})_{n}b = ∑_{i≥0} a_{−1−i}(a_{n+i}^{k}b) + ∑_{i≥0} a_{n−1−i}^{k}(a_{i}b)

which is a special case of the vertex algebra identity. This proves lemma 3.1.

Corollary 3.2. Suppose that a is an element of a vertex algebra V such that a_{i}a = 0 for i ≥ −1. Then for each k ≥ 0 the operator a_{0}^{k} is divisible by k!.

Proof. We show that a_{0}^{k} is divisible by k! by induction on k, which is obvious for k = 0, 1. As a_{−1}a = 0 the left hand side of lemma 3.1 vanishes for k ≥ 2. As a_{i}a = 0 for i ≥ 0 all the operators a_{i} commute with each other because of the formula

[a_{i}, b_{j}] = ∑_{k≥0} \binom{i}{k}(a_{k}b)_{i+j−k},

which together with the induction hypothesis implies that the sum of all terms on the right hand side of lemma 3.1 other than a_{0}^{k} is divisible by k!. Hence a_{0}^{k} is also divisible by k!. This proves lemma 3.2.

Corollary 3.3. Suppose V is the vertex algebra (over Z) of a double cover L̂ of an even integral lattice L. If α ∈ L̂ has positive norm and a is the element e^{α} of V, then all coefficients of

exp(xa_{0}) = ∑_{k} x^{k}a_{0}^{k}/k!

map V into V.

Proof. For any α, β ∈ L we have e_{i}^{α}e^{β} = 0 for i + (α, β) ≥ 0. In particular e_{i}^{α}e^{α} = 0 for i ≥ −1 if (α, α) > 0. Corollary 3.3 now follows from corollary 3.2.

Corollary 3.3 will allow us to lift elements in the root spaces of norm 2 vectors. In the rest of this section we show how to lift elements in the root spaces of norm 0 vectors.

Lemma 3.4. Suppose that the the elements Γ_{i} for i ∈ Z are independent formal variables. If I = (i_{1}, …, i_{m}) is a finite sequence of integers then define Γ_{I} to be Γ_{i_{1}} ⋯ Γ_{i_{m}} and define Σ(I) to be i_{1} + ⋯ + i_{m} and define l(I) to be m. Then all coefficient of the power series

E = exp ( ∑_{m,n>0} ∑_{I∈Z^{m}} ∑_{{J∈Z^{n} \atop Σ(J)=−Σ(I)>0}} \frac{Σ(I)}{mn}Γ_{I}Γ_{J})

are integers.

Proof. If I is any finite sequence of integers then I^{k} means the obvious concatenation of k copies of the sequence I. We will call a pair of finite sequences primitive if it is not of the form (I^{k}, J^{k}) for some k ≥ 2. Any pair of finite sequences can be written uniquely as (I^{k}, J^{k}) for some primitive pair (I, J), which we call the primitive core of the pair (I^{k}, J^{k}). Consider the group Z × Z acting on pairs (I, J) by the first Z acting as cyclic permutations of the elements of I and the second Z acting as cyclic permutations of the elements of J. We group the pairs (I, J) indexing the terms of the sum in the exponent of E into equivalence classes, where we say two pairs of sequences are equivalent if their primitive cores are conjugate under Z × Z. For any primitive element (I, J) with Σ(I) + Σ(J) = 0 we will show that all coefficients of the exponential of the sum of all terms in the equivalence class of (I, J) are integral. This will show that the coefficients of E are integral because E is an infinite product over the set of orbits of primitive elements of expressions like this. Let m′ and n′ be the number of orbits of I and J under the cyclic action of Z. Then the exponential of the sum of the terms equivalent to (I, J) is

exp ( ∑_{k>0} \frac{Σ(I^{k})}{l(I^{k})l(J^{k})}(number of orbits of (I, J) under Z × Z)Γ_{I^{k}}Γ_{J^{k}})

= exp ( ∑_{k>0} \frac{kΣ(I)m′n′}{kmkn} Γ_{I^{k}}Γ_{J^{k}})

= exp (\frac{Σ(I)m′n′}{mn} ∑_{k>0} Γ_{I}^{k}Γ_{J}^{k}/k)

= exp ( − \frac{Σ(I)m′n′}{mn} log(1 − Γ_{I}Γ_{J}))
=(1 − Γ_{I}Γ_{J})^{−Σ(I)m′n′/mn}

The number Σ(I) is divisible by m/m′ as I is the concatenation of m/m′ identical sequences and Σ(I) is the sum of the elements of I, and similarly Σ(I) = −Σ(J) is divisible by n/n′. Moreover m/m′ and n/n′ are coprime as (I, J) is primitive. Hence Σ(I)m′n′/mn is an integer because it is equal to Σ(I) divided by two coprime factors m/m′ and n/n′ of Σ(I). This implies that (1 − Γ_{I}Γ_{J})^{−Σ(I)m′n′/mn} has integral coefficients, and so E does as well as it is an infinite product of expressions like this. This proves lemma 3.4.

Theorem 3.5. Let V be the vertex algebra (over the integers) of the double cover of an even lattice L. Define D^{∗}V to be the sum of the spaces D^{(i)}V for i ≥ 1. We recall from [B86] that V/D^{∗}V has a natural Lie algebra structure, with the bracket defined by [u, v] = u_{0}v. Moreover this Lie algebra acts on V preserving the vertex algebra structure of V. Suppose that α, γ ∈ L with α orthogonal to the norm 0 vector γ. Then

exp ( ∑_{i>0} \frac{x^{i}}{i} (α(1)e^{iγ})_{0})

is a lifting of (α(1)e^{γ})_{0} in the universal enveloping algebra of (V/D^{∗}V) ⊗ Q all of whose coefficients map V to V.

Proof. It is obvious that the element is group-like as it is the exponential of a primitive element, so the only problem is to show that it preserves the integral form V.

This element is also an automorphism of the vertex algebra (V ⊗ Q)[[x]] as it is the exponential of a derivation of this vertex algebra. Hence to show its coefficients preserve the integral form of V it is sufficient to show that it maps each of the generators e^{β}, β ∈ L, of the vertex algebra V into V [[x]].

Define the operators Γ_{i} by

e^{γ}(z) = ∑_{i} Γ_{i}z^{i}.

Recall the following formulas from [B86].

α(1)(z) = ∑_{j} α(1)_{−j}z^{j−1} = ∑_{j} α(j)z^{j−1}

All coefficients of e^{γ} commute with everything in sight as γ is orthogonal to γ and α. So (α(1)e^{iγ})_{0} is equal to the coefficient of z^{−1} in

α(1)(z)(e^{γ}(z))^{i} = ∑_{j} α(j)z^{j−1} ∑_{I∈Z^{i}} Γ_{I}z^{Σ(I)}

Therefore ∑_{i>0} \frac{x^{i}}{i} (α(1)e^{iγ})_{0} is equal to A^{+} + A^{0} + A^{−} where

A^{+} = ∑_{m>0} ∑_{j>0} α(j) ∑_{{I∈Z^{m} \atop Σ(I)=−j}} \frac{x^{m}}{m} Γ_{I}

A^{0} = ∑_{m>0} ∑_{j=0} α(j) ∑_{{I∈Z^{m} \atop Σ(I)=−j}} \frac{x^{m}}{m} Γ_{I}

A^{−} = ∑_{m>0} ∑_{j<0} α(j) ∑_{{I∈Z^{m} \atop Σ(I)=−j}} \frac{x^{m}}{m} Γ_{I}

We would like to pull out a factor of exp(A^{−}) from exp(A^{+} + A^{0} + A^{−}) because exp(A^{−})(e^{β}) = e^{β}, but we have to be careful when doing this because α(i) does not commute with α(−i) if i ≠ 0.

We can easily evaluate [A^{+}, A^{−}]/2 using the fact that [α(i), α(j)] = j(α, α) if i+j = 0 and 0 otherwise, and we find that

[A^{+}, A^{−}]/2 = \frac{1}{2}(α, α) ∑_{m,n>0} ∑_{k<0} ∑_{{I∈Z^{m} \atop Σ(I)=k}} ∑_{{J∈Z^{n} \atop Σ(J)=−k}} \frac{k}{mn}Γ_{I}Γ_{J}.

In particular this verifies the fact used below that [A^{+}, A^{−}] commutes with A^{+} and A^{−}. The exponential of this is the expression in lemma 3.4 raised to the power of (α, α)/2, and therefore has integral coefficients by lemma 3.4 and because (α, α) has even norm. So exp([A^{+}, A^{−}]/2) maps V into V [[x]].

Recall the formula

exp(A^{+} + A^{−}) = exp([A^{−}, A^{+}]/2) exp(A^{+}) exp(A^{−})

which is valid because [A^{+}, A^{−}] commutes with A^{+} and A^{−}. (This is essentially just the first few terms of the Baker-Campbell-Hausdorff formula.) Now we look at

exp(A^{+} + A^{0} + A^{−})e^{β} = exp([A^{−}, A^{+}]/2) exp(A^{+}) exp(A^{0}) exp(A^{−})e^{β}

It is obvious that exp(A^{−})(e^{β}) = e^{β}, and we have checked above that all coefficients of exp([A^{−}, A^{+}]/2) map V to V. Hence to complete the proof of theorem 3.5 it is sufficient to prove that all coefficients of exp(A^{0}) and exp(A^{+}) map V into V.

We check that exp(A^{0}) maps V to V [[x]]. The follows because A^{0} is an infinite sum of expressions like

∑_{n>0} α(0)Γ_{J^{n}}x^{n}/n = −α(0) log(1 − xΓ_{J})

where J is a primitive sequence (in other words a sequence that cannot be written in the form I^{m} for some m > 1). Hence exp(A^{0}) is an infinite product of terms of the form (1 − xΓ_{J})^{−α(0)}, which map V [[x]] to itself because α(0) has integral eigenvalue (α, β) on the subspace of V of degree β ∈ L. This shows that all coefficients of exp(A^{0}) map V into V.

Finally we have to show that all coefficients of exp(A^{+}) map V into V. As usual we divide the sum over elements I in A^{+} into classes consisting of powers of conjugates of primitive elements I. We see that A^{+} is a sum over all orbits of primitive elements I with Σ(I) > 0 of expressions like

l(I) ∑_{k>0} α(Σ(I^{k}))\frac{x^{l(I^{k})}}{l(I^{k})} Γ_{I^{k}} = ∑_{k>0} α(kΣ(I))\frac{x^{kl(I)}}{k} Γ_{I}^{k}.

(The factor at the front is the number of conjugates of I under the cyclic action of Z, which is equal to l(I) because I is primitive.) So it is sufficient to show that the exponential of this expression has integral coefficients. Let y_{1}, y_{2}, …, be a countable number of independent variables, and identify α(k) with the symmetric function ∑ y_{i}^{k} of the y’s for k > 0. (See [M chapter 1].) Then

exp ( ∑_{k>0} α(kΣ(I))\frac{x^{kl(I)}}{k} Γ_{I}^{k})

= exp ( ∑_{i} ∑_{k>0} y_{i}^{kΣ(I)} \frac{x^{kl(I)}}{k} Γ_{I}^{k})
= exp ( ∑_{i} − log(1 − y_{i}^{Σ(I)}x^{l(I)}Γ_{I}))

= ∏_{i} \frac{1}{1 − y_{i}^{Σ(I)}x^{l(I)}Γ_{I}} .

So we see that

exp(A^{+}) = ∏_{I} ∏_{i} \frac{1}{1 − y_{i}^{Σ(I)}x^{l(I)}Γ_{I}}

where the product over I is a product over all orbits of primitive sequences I with Σ(I) > 0 under the cyclic action of Z. The last line is a power series in the elements Γ_{I} and x whose coefficients are symmetric functions in the y’s, and hence are polynomials with integral coefficients in the complete symmetric functions of the y’s. So we have to show that each complete symmetric function, considered as a polynomial in the α’s with rational (not necessarily integral!) coefficients maps V to V. The complete symmetric functions are the polynomials e^{−α}D^{(n)}(e^{α}) considered as elements of the ring V underlying the vertex algebra V (as follows from [M, Chapter 1, 2.10]). By definition of the integral form V these polynomials map V to itself.

This proves theorem 3.5.

4. The fake monster smooth Hopf algebra.

We recall the construction of the fake monster Lie algebra [B90]. It is the Lie algebra of physical states of the vertex algebra of the double cover of the lattice II_{25,1}. This Lie algebra has an integral form m consisting of the elements represented by elements of the integral form of the vertex algebra. We recall some properties of m:

1 m is graded by the lattice II_{25,1}, and the piece m_{α} of degree α ∈ II_{25,1} has dimension p_{24}(1 − α^{2}/2) if α ≠ 0 and 26 if α = 0, where p_{24}(n) is the number of partitions of n into parts of 24 colors.

2 m has an involution ω lifting the involution −1 of II_{25,1}.

3 m has a symmetric invariant integer valued bilinear form (,), and the pairing between

m_{α} ⊗ Q and m_{−α} ⊗ Q is nonsingular. 4 m ⊗ Q is a generalized Kac-Moody algebra. The simple roots are given by the norm

2 vectors r with (r, ρ) = −1, together with all positive multiples of ρ with multiplicity 24, where ρ is a primitive norm 0 vector of II_{25,1} such that ρ^{⊥}/ρ is isomorphic to the Leech lattice. (This follows from [B90. theorem 1].) We define U^{+}(m) to be the Z-subalgebra of the universal enveloping algebra U(m⊗Q)

generated by the coefficients of the liftings of elements in root spaces of the simple roots and their negatives constructed in corollary 3.3 and theorem 3.5. (Note that all simple roots of m have norms 2 or 0 so we can always apply one of these two types of liftings.)

Theorem 4.1. There is a II_{25,1}-graded Hopf algebra U^{+}(m) over Z with the following properties.

1 U^{+}(m) has a structural basis over Z.

2 The primitive elements of U^{+}(m) are an integral form of the fake monster Lie algebra

m. 3 For every norm 2 vector of II_{25,1}, U^{+}(m) contains the usual (Kostant) integral form

of the universal enveloping algebra of the corresponding sl_{2}(Z).

Proof. The algebra U^{+}(m) is a Z-Hopf subalgebra of U(m ⊗ Q) as it is generated by coalgebras. Also U^{+}(m) is obviously torsion free as it is contained in a rational vector space. It is easy to check directly that the degree zero primitive elements of U^{+}(m) are just the degree 0 elements of m and therefore form a free Z module (of rank 26). If α is any nonzero element of II_{25,1} having nonzero inner product with some element β ∈ II_{25,1}, then (α, β)u ∈ m_{α} for any primitive element u ∈ U^{+}(m) of degree α, because U^{+}(m) maps m, and hence β, to m by corollary 3.3 and theorem 3.5. This shows that all root spaces of primitive elements of U^{+}(m) are free Z-modules. We can now apply theorem 2.15 to see that U^{+}(m) is a Hopf algebra with a structural basis. The primitive elements of U^{+}(m) form an integral form of the fake monster Lie algebra, because the fake monster Lie algebra over the rationals is generated by the root spaces of simple roots and their negatives, and all simple roots have norms 2 or 0. This proves theorem 4.1.

5. A smooth Hopf algebra for the Virasoro algebra.

In this section we show that there is a Hopf algebra over Z with a structural basis whose primitive elements form the natural integral form of the Virasoro algebra (theorem 5.7). In other words, there is a formal group law over Z corresponding to the Virasoro algebra. Moreover this Hopf algebra acts on the integral form of the vertex algebra of any even self dual lattice.

Let R be a commutative ring. We write Hom(R, R) for the ring of homomorphisms of the abelian group R to itself, and Der(R) for the Lie algebra of derivations of the ring R, and U(Der(R)) for the universal enveloping algebra of Der(R) over Z. Consider the group of all element a = ∑_{i} a_{i}ϵ^{i} ∈ Hom(R, R)[[ϵ]] with a_{0} = 1 that induce automorphisms of the Z[[ϵ]] algebra R[[ϵ]]. We can think of the elements of this group informally as “infinitesimal curves in the group of automorphisms of Spec(R)”. We will call a derivation of R liftable if it is of the form a_{1} for some a as above.

Lemma 5.1. Let R be any commutative algebra with no Z-torsion. Then any a ∈ Hom(R, R)[[ϵ]] with a(0) = 1 that is an automorphism of R[[ϵ]] is the image of a unique group-like element G_{a} of U(Der(R) ⊗ Q)[[ϵ]] with G_{a}(0) = 1 under the natural map from U(Der(R) ⊗ Q) to Hom(R, R) ⊗ Q.

Proof. We note that log(a) is a well defined element of (Hom(R, R) ⊗ Q)[[ϵ]] as a = 1 + O(ϵ) ∈ Hom(R, R)[[ϵ]]. As a is a ring homomorphism it follows that that log(a) is a derivation of (R ⊗ Q)[[ϵ]] and is therefore an element of (Der(R) ⊗ Q)[[ϵ]]. Now we consider log(a) to be an element of the universal enveloping algebra U(Der(R) ⊗ Q)[[ϵ]] and we define G_{a} ∈ U(Der(R) ⊗ Q)[[ϵ]] by

G_{a} = exp(log(a)),

where the log is computed in (Hom(R, R) ⊗ Q)[[ϵ]] and the exponential is computed in U(Der(R) ⊗ Q)[[ϵ]]. It is obvious that the action of G_{a} on R[[ϵ]] is the same as that of a. Also G_{a} is group-like because it is the exponential of a primitive element. It is easy to check that G_{a} is the unique group-like lifting of a with G_{a}(0) = 1, because the log of a group-like element must be primitive and must therefore be the same as log(a). This proves lemma 5.1.

Corollary 5.2. Suppose that R is a commutative ring with no Z torsion such that Der(R) is a free Z-module. Define U^{+}(Der(R)) to be the subalgebra of U(Der(R ⊗ Q)) generated by all the coefficients of all group-like elements of U(Der(R) ⊗ Q)[[ϵ]] that have constant coefficient 1 and map R[[ϵ]] to R[[ϵ]]. Then U^{+}(Der(R)) is a Hopf algebra over Z with a structural basis, and its primitive elements are the liftable primitive elements of Der(R).

Proof. Applying theorem 2.15 shows that U^{+}(Der(R)) is a Hopf algebra over Z with a structural basis. By lemma 5.1 the space of primitive elements of U^{+}(Der(R)) is the same as the space of liftable primitive elements of Der(R). This proves corollary 5.2.

Example 5.3 Suppose we take R to be the algebra Z[x][x^{−1}] of Laurent polynomials. Then Der(R) = W itt is the Witt algebra over Z, which is spanned by the elements L_{m} = −x^{m+1} \frac{d}{dx} for m ∈ Z. All elements L_{m} are liftable; for example, we can use the automorphism of R[[ϵ]] taking x to x − ϵx^{m+1} to show that L_{m} is liftable. Therefore the Hopf algebra U^{+}(W itt) is a Hopf algebra over Z with a structural basis, whose primitive elements are exactly the elements of the Witt algebra.

Lemma 5.4. Let N be an integer. Let R_{N} be the representation of W itt with a basis of elements e_{n}, n ∈ Z, with the action given by L_{m}(e_{n}) = (Nm + n)e_{m+n}. Then the action of W itt on R_{N} can be extended to an action of U^{+}(W itt) on R_{N}.

Proof. For N = −1 the module R_{N} is the module of first order differential operators on R = Z[x][x^{−1}], so the automorphisms G_{a} extend to R_{N}[[ϵ]]. For other negative values of N the R-module R_{N} is a tensor product of −N copies of the module R_{−1}, and R_{N} for N positive is the dual of R_{N} for N negative. Therefore U^{+}(W itt) extends to these modules as well. This proves lemma 5.4.

We define W itt_{≥n} for n = −1, 1 to be the subalgebra of W itt spanned by L_{i} for i ≥ n. Let V be the vertex algebra of some even lattice. It contains elements e^{α} for α ∈ L, so there are operators e_{i}^{α} on V for i ∈ Z. The algebra W itt_{≥−1} also acts naturally on V.

Lemma 5.5. Put N = α^{2}/2 − 1 and e_{j} = e_{α^{2}/2−1−j}^{α} for j ∈ Z and let R_{N} be the space with the elements e_{j} as a basis. Define an action of the algebra U^{+}(W itt_{≥−1}) on R_{N} ⊗U^{+}(W itt_{≥−1}) using the action on R_{N} as in lemma 5.4 and the action on U^{+}(W itt_{≥−1}) by left multiplication and the coalgebra structure of U^{+}(W itt_{≥−1}). If u ∈ U^{+}(W itt_{≥−1}) then ue_{m}^{α} = ∑_{i} e_{i}^{α}u_{i} as operators on V, where ∑_{i} e_{i}^{α} ⊗ u_{i} = u(e_{m}^{α} ⊗ 1) is the image of e_{m}^{α} ⊗ 1 under the action of u on R_{N} ⊗ U^{+}(W itt_{≥−1}). In particular if U^{+}(W itt_{≥−1}) maps some element v ∈ V to V then it also maps e_{i}^{α}(v) to V for any α ∈ L and i ∈ Z.

Proof. If u is of the form L_{i} for i ≥ −1 this can be proved as follows. A standard vertex algebra calculation shows that

[L_{i}, e_{j}^{α}] = ((i + 1)(α^{2}/2 − 1) − j)e_{i+j}^{α}.

This shows that lemma 5.5 is true when u ∈ W itt_{≥−1}. If lemma 5.5 is true for two elements u, u′ of U^{+}(W itt_{≥−1}) then it is true for their product. If it is true for some nonzero integral multiple of u ∈ U^{+}(W itt_{≥−1}) then it is true for u because V is torsion-free. To finish the proof we observe that the algebra U^{+}(W itt_{≥−1}) is generated up to torsion by the elements L_{m} for m ≥ −1. This proves lemma 5.5.

Lemma 5.6. Suppose that V is the vertex algebra of the double cover of some even lattice, with the standard action of W itt_{≥−1} on V ⊗ Q ([B86]). This action extends to an action of U^{+}(W itt_{≥−1}) on V.

Proof. The vertex algebra V is generated from the element 1 by the actions of the operators e_{n}^{α} for n ∈ Z and α ∈ L. Define F^{n}(V) by defining F^{0}(V) to be the space spanned by 1, and defining F^{n+1}(V) to be the space spanned by the actions of operators of the form e_{m}^{α} on F^{n}(V). Then V is the union of the spaces F^{n}(V), so it is sufficient to prove that each space F^{n}(V) is preserved by the action of U^{+}(W itt_{≥−1}). We will prove this by induction on n. For n = 0 is is trivial because L_{n}(1) = 0 for n ≥ −1. If it is true for n, then it follows immediately from lemma 5.5 that it is true for n + 1. This proves lemma 5.6.

The algebra U^{+}(W itt_{≥−1}) can be Z-graded in such a way that L_{m} has degree m. This follows easily from the fact that we can find graded liftings of the elements L_{m}. We define U^{+}(W itt_{≥1}) to be the subalgebra generated by the coefficients of graded liftings of the elements L_{m} for m ≥ 1. It is easy to check that this is a Hopf algebra with a structural basis whose Lie algebra of primitive vectors has a basis consisting of the elements L_{m} for m ≥ 1. It is Z_{≥0}-graded, with all graded pieces being finite dimensional; in fact the piece of degree n ∈ Z has dimension p(n) where p is the partition function.

We let U(W itt_{≥1}) be the universal enveloping algebra of the Lie algebra W itt_{≥1}. It is Z_{≥0}-graded in the obvious way and is a subalgebra of U^{+}(W itt_{≥1}).

In the rest of this section we construct an integral form with a structural basis for the universal enveloping algebra of the Virasoro algebra over Z. This result is not used elsewhere in this paper. We recall that the Virasoro algebra V ir is a central extension of W itt and is spanned by elements L_{i} for i ∈ Z and an element c/2 in the center, with

[L_{m}, L_{n}] = (m − n)L_{m+n} + \binom{m + 1}{3}\frac{c}{2}.

We identify W itt_{≥−1} with the subalgebra of V ir spanned by L_{m} for m ≥ −1. The Virasoro algebra V ir has an automorphism ω of order 2 defined by ω(L_{m}) = −L_{−m}, ω(c) = −c, and ω extends to an automorphism of the universal enveloping algebra U(V ir ⊗ Q). We define U^{+}(V ir) to be the subalgebra of U(V ir ⊗ Q) generated by U^{+}(W itt_{≥−1}) and ω(U^{+}(W itt_{≥−1})).

For any even integral lattice L there is a double cover L̂, unique up to non-unique isomorphism, such that e^{a}e^{b} = (−1)^{(a,b)}e^{b}e^{a}. We let VL̂ be the (integral form of the) vertex algebra of L̂. This is L-graded, and has a self dual bilinear form on it (more precisely, each piece of given L-degree α and given eigenvalue under L_{0} is finite dimensional and dual to the piece of degree −α), and if L is self dual then VL̂ has a conformal vector generating an action of the Virasoro algebra.

Theorem 5.7. The subalgebra U^{+}(V ir) of U(V ir ⊗ Q) has the following properties: 1 U^{+}(V ir) is a Z-Hopf algebra with a structural basis. 2 The Lie algebra of primitive elements of U^{+}(V ir) has a basis consisting of the elements

L_{n} for n ∈ Z and the element c/2. 3 U^{+}(V ir) maps the vertex algebra (over Z) of any even self dual lattice to itself.

Proof. We first construct the action on the vertex algebra V of an even self dual lattice. We have an action of U^{+}(W itt_{≥−1}) on V by lemma 5.6. The vertex algebra of an even self dual lattice is also self dual under its natural bilinear form, so the adjoint of any linear operator on V is also a linear operator on V. The adjoint of L_{m} is L_{−m}, so the adjoint U^{+}(W itt_{≤1}) = ω(U^{+}(W itt_{≥−1})) also maps V to itself. As these two algebras generate U^{+}(V ir) this proves that U^{+}(V ir) acts on V.

Next we find the Lie algebra P of primitive elements of U^{+}(V ir). The elements L_{m} for m ≥ −1 are obviously in U^{+}(V ir) because they are in U^{+}(W itt_{≥−1}), and similarly L_{m} for m ≤ 1 is in P. The element c/2 is in P because [L_{2}, L_{−2}] = 4L_{0} + c/2. So P contains the basis described in theorem 5.7, and we have to prove that P contains no elements other than linear combinations of these. As U^{+}(V ir) and hence P are both Z-graded with L_{m} having degree m, it is sufficient to show that the degree m piece of P is spanned by L_{m} if m ≠ 0 and by L_{0} and c/2 if m = 0. For m ≠ 0 this is easy to check as we just map the Virasoro algebra to the Witt algebra and use example 5.3. The case m = 0 is harder and we will use the actions on vertex algebras of even self dual lattices L constructed above. If L is such a lattice then c acts on V as multiplication by dim(L), and L_{0} has eigenspaces with eigenvalue any given positive integer (at least if L has positive dimension). So if xL_{0} +yc/2 is in P for some x, y ∈ Q then xm+yn/2 is an integer whenever m is a positive integer and n is the dimension of a nonzero even self dual lattice. As we can find such lattices for any positive even integer n, this implies that x and y are both integers. Hence the degree 0 piece of P is spanned by L_{0} and c/2. This completes the proof that P is spanned by L_{m}, m ∈ Z, and c/2.

Finally we have to show that U^{+}(V ir) has a structural basis. We know that all the elements L_{m} for m ≥ −1 are liftable in U^{+}(W itt_{≥−1}), so they are also liftable in U^{+}(V ir). Similarly the elements L_{m} for m ≤ 1 are liftable in ω(U^{+}(W itt_{≥−1})) and therefore in U^{+}(V ir). It is trivial to check that c is liftable as it acts as multiplication by some integer, so every primitive element of U^{+}(V ir) is liftable by lemma 2.5 and part 2 of theorem 5.7. Also U^{+}(V ir) is generated by the coefficients of group-like elements because this is true for U^{+}(W itt_{≥−1}) and its conjugate under ω. The fact that U^{+}(V ir) has a structural basis now follows from theorem 2.15. This completes the proof of theorem 5.7.

Example. If L is an even lattice of odd dimension then L_{−2}(1) is not in the vertex algebra of L, because [L_{2}, L_{−2}] = 4L_{0} +dim(L)/2. So part 3 of theorem 5.7 is false without the assumption that L is self dual.

6. The no-ghost theorem over Z.

The no-ghost theorem of Goddard and Thorn [G-T] implies that over the reals, the contravariant form restricted to the degree β ≠ 0 piece of the fake monster Lie algebra m is positive definite and in particular nonsingular. We give a refinement of this to the integral form of the degree β piece m_{β}, showing that the any prime dividing the discriminant of the quadratic form on m_{β} also divides the vector β. In particular if β is a primitive vector then m_{β} is a self dual positive definite integral lattice. I do not know whether or not the discriminant can be divisible by p when β is divisible by p.

If λ = 1^{i_{1}}2^{i_{2}} ⋯ is a partition with i_{1} 1’s, i_{2} 2’s, and so on, then we define P (λ) to be the integer 1^{i_{1}}2^{i_{2}} ⋯ and F (λ) to be i_{1}!i_{2}! ⋯ and |λ| to be 1i_{1} + 2i_{2} + ⋯ and l(λ) to be i_{1} + i_{2} + · · ·. We also define p(n) to be the number of partitions of n.

Lemma 6.1. Suppose that n is an integer. Then

∏_{|λ|=n} P (λ) = ∏_{|λ|=n} F (λ).

Proof. We will show that both sides are equal to

∏_{i>0} i^{Σ_{j>0}p(n−ij)}.

The left hand side ∏_{|λ|=n} P (λ) is equal to ∏_{i>0} i^{n(i)} where n(i) is the number of times i occurs in some partition of n, counted with multiplicities. This number n(i) is equal to ∑_{j} n(i, j) where n(i, j) is the number of times that i occurs at least j times in a partition of n. But n(i, j) is equal to p(n − ij) because any partition of n in which i occurs at least j times can be obtained uniquely from a partition of n − ij by adding j copies of i. So the left hand side ∏_{|λ|=n} P (λ) is equal to ∏_{i>0} i^{Σ_{j>0}p(n−ij)}.

On the other hand the right hand side ∏_{|λ|=n} F (λ) is equal to ∏_{i>0} i^{m(i)} where m(i) is the number of times that there is a partition of n with some number occurring at least i times (counting a partition several times if it has more than one number occurring at least i times). But m(i) is equal to ∑_{j} n(j, i) = ∑_{j} p(n − ji). This shows that the right hand side ∏_{|λ|=n} F (λ) is also equal to ∏_{i>0} i^{Σ_{j>0}p(n−ij)}. This proves lemma 6.1.

Lemma 6.2. The submodule U(W itt_{≥1})_{n} has index ∏_{|λ|=n} F (λ) inside U^{+}(W itt_{≥1})_{n}.

Proof. Choose a graded lifting 1 + a_{i,1}x + a_{i,2}x^{2} + … of L_{i} for each i > 0. Then the elements a_{1,i_{1}}a_{2,i_{2}} ⋯ for 1i_{1} + 2i_{2} + ⋯ = n form a base for U^{+}(W itt_{≥1}), and the elements i_{1}!a_{1,i_{1}}i_{2}!a_{2,i_{2}} ⋯ for 1i_{1} + 2i_{2} + ⋯ = n form a basis for U(W itt_{≥1})_{n}. Therefore the index of U(W itt_{≥1})_{n} in U^{+}(W itt_{≥1})_{n} is

∏_{i_{1}+2i_{2}+···=n} i_{1}!i_{2}! ⋯ = ∏_{|λ|=n} F (λ).

This proves lemma 6.2.

Lemma 6.3. Let γ be a norm 0 vector of II_{25,1}. Suppose that W is the graded space of all elements generated by the action of the elements e^{−γ}D^{(i)}(e^{γ}) on e^{β}, so that W is acted on by the smooth integral form U = U^{+}(W itt_{>0}). Then the graded dual W [1/(β, γ)]^{∗} of W [1/(β, γ)] is a free U[1/(β, γ)]-module on one generator.

Proof. We define U_{n} and W_{n} to be the degree n pieces of U and W. Let w_{μ} be the basis of elements γ(1)^{j_{1}}γ_{2}^{j_{2}} ⋯ e^{β} for W_{n} ⊗ Q parameterized by partitions μ = 1^{j_{1}}2^{j_{2}} ⋯ of n. The Z module W′_{n} spanned by the w_{λ}’s is not W_{n} but has index ∏_{|μ|=n} F (μ) in it. We let the elements L_{λ} = L_{1}^{i_{1}}L_{2}^{i_{2}} ⋯ be the basis for the space U_{n} ⊗ Q indexed by partitions λ = 1^{i_{1}}2^{i_{2}} · · ·. The Z module U′_{n} spanned by the L_{λ}’s is not U_{n} but has index ∏_{|λ|=n} F (λ) in it by lemma 6.2. We define m_{λ,μ} for |λ| = |μ| by L_{λ}(w_{μ}) = m_{λ,μ}e^{β}. We will show that the determinant of the p(n) by p(n) matrix (m_{λ,μ}) is

∏_{|λ|=n}(β, γ)^{l(λ)}P (λ)F (λ)

where l(λ) is the number of elements of the partition λ. We order the partitions by λ > μ if λ is the partition λ_{1} + λ_{2} + ⋯ with λ_{1} ≥ λ_{2} ≥ ⋯, μ is the partition μ_{1} + μ_{2} + ⋯ with μ_{1} ≥ μ_{2} ≥ ⋯, and λ_{1} = μ_{1}, …, λ_{k−1} = μ_{k−1}, λ_{k} > μ_{k} for some k. Then the matrix entry m_{λ,μ} is 0 if λ > μ, and is equal to

P (λ)F (Λ)(β, γ)^{l(λ)}

if λ = μ. We can see this by repeatedly using the relation

⋯ L_{i_{2}}L_{i_{1}}γ(j_{1})γ(j_{2}) ⋯ e^{β}
= \{ \begin{matrix}(number of j’s equal to j_{1})(β, γ)j_{1} ⋯ L_{i_{2}}γ(i_{2}) ⋯ e^{β} & if i_{1} = j_{1} \\ 0 & if i_{1} > j_{1} \end{matrix}

for j_{1} ≥ j_{2} ≥ ⋯, which in turn follows from the identities [L_{i}, γ(j)] = jγ(j − i), L_{i}(e^{β}) = 0 = γ(−i)(e^{β}) for i > 0, and γ_{0}e^{β} = (β, γ) As the matrix (m_{λ,μ}) is triangular its determinant is given by the product of the diagonal entries m_{λ,λ} = P (λ)F (Λ)(β, γ)^{l(λ)}.

Now we work out the index of U^{+}(W itt_{>0})(e^{β^{∗}})_{n} in W_{n}^{∗} where e^{β∗} is the basis element of W_{0}^{∗} dual to e^{β} ∈ W_{0}. This index is equal to

\frac{(Index of U′_{n} in W′_{n})}{(Index of U′_{n} in U_{n})(Index of W′_{n} in W_{n})}

The numerator of this expression is equal to the determinant of the matrix (m_{λ,μ}), and we calculated this earlier. Substituting in the known values we find the index of U^{+}(W itt_{>0})(e^{β∗})_{n} in W_{n}^{∗} is

\frac{∏_{|λ|=n} P (λ)F (λ)(β, γ)^{l(λ)}}{(∏_{|λ|=n} F (λ))(∏_{|μ|=n} F (μ))}

Applying lemma 6.1 we see that this is equal to ∏_{|λ|=n}(β, γ)^{l(λ)}. This is a unit in Z[1/(β, γ)], so over the ring Z[1/(β, γ)] the map from U^{+}(W itt_{>0}) to W^{∗} is an isomorphism. This proves lemma 6.3.

Fix a norm 0 vector γ ∈ L. We recall that the transverse space is the subspace of elements v ∈ V such that L_{i}(v) = 0 for i > 0, L_{0}(v) = v, and γ(i)(v) = 0 for i < 0. It is easy to check that the transverse space T_{β} of degree β with (β, γ) ≠ 0 is positive definite, and the no-ghost theorem [G-T] works by showing that the natural map from T_{β} to the space of physical states (modulo null vectors) of degree β is an isomorphism.

Lemma 6.4. The transverse space T_{β} of degree β ∈ II_{1,1} has determinant dividing a power of (β, γ).

Proof. For any vector v of V there is a U^{+}(W itt_{>0}) equivariant map from W^{∗} to V taking 1 to v over Z[1/(β, γ)] by lemma 6.3. Dualizing we see that this means there is a vector in W ⊗ V of the form v ⊗ 1+ (terms involving some γ(i)) which is fixed by U^{+}(W itt_{>0}) and hence also fixed by W itt_{>0}. This gives a map from V to the transverse space which is an isomorphism over Z[1/(β, γ)]. This isomorphism preserves the bilinear form because γ has norm 0 so all the terms involving γ(i) have zero inner product with all other terms. Hence the transverse space is self dual over Z[1/(γ, β)]. This proves lemma 6.4.

The next theorem is an extension of the no-ghost theorem from rational vector spaces to modules over Z. Recall that the usual no-ghost theorem [G-T] says that if α ∈ II_{25,1} is nonzero then the space of physical states (over Q) of degree α is positive definite and spanned by the transverse space, which has dimension p_{24}(1 − α^{2}/2). This describes the space of physical states as a rational vector space, but it also has a natural lattice inside it and we can ask about the structure of this lattice.

Theorem 6.5. (The no-ghost theorem over Z). Suppose α is a nonzero vector of II_{25,1} which is n times a primitive vector. Then the discriminant of the space of physical states of degree α divides a power of n. In particular the space of physical states is a self dual lattice if β is primitive.

Proof. For every prime p coprime to n we can find a norm zero vector γ with (γ, β) coprime to p, because the norm 0 vectors of Leech type span II_{25,1}. By lemma 6.4 this implies that the discriminant of the space of physical states of degree β is coprime to p. Hence this discriminant divides a power of n. This proves theorem 6.5.

The proof of this theorem implies that all spaces of physical states of primitive vectors of the same norm are isomorphic over all p-adic fields, in other words in the same genus. It is certainly not always true that they are isomorphic over the integers. For example, for a norm 0 vector the space of physical states is isomorphic to the corresponding Niemeier lattice, and not all Niemeier lattices are isomorphic.

7. An application to modular moonshine.

In this section m will stand for the monster Lie algebra over Z[1/2] (see [B-R], [B98]) rather than the fake monster Lie algebra.

We write Z_{p} for the ring of p-adic numbers. The paper [B98] showed that Ryba’s modular moonshine conjectures [R] for primes p ≥ 13 were true provided the following assumption was true:

Assumption. If m < p then the degree (m, n) piece of m⊗Z_{p} is self dual under the natural bilinear form and isomorphic to V_{mn} ⊗ Z_{p} as a Z_{p} module acted on by the monster.

In this section we will prove this assumption, thus completing the proof of the modular moonshine conjectures for p ≥ 13. In fact we will prove the following slightly stronger theorem:

Theorem 7.1. If m and n are not both divisible by p then the degree (m, n) piece of m ⊗ Z_{p} is self dual under the natural bilinear form and isomorphic to V_{mn} ⊗ Z_{p} as a Z_{p} module acted on by the monster.

Proof. If one of m or n is not divisible by p then we can find a norm 0 vector γ with (γ, β) coprime to p. Theorem 7.1 then follows from the integral no-ghost theorem 6.5, because the discriminant of m_{(m,n)} is not divisible by p and is therefore a unit in Z_{p}. Note that the isomorphism given by theorem 6.5 preserves the action of the monster. This proves theorem 7.1.

8. Open problems.

1 The integral form for the fake monster Lie algebra constructed in section 4 is probably not the best possible one. Is there an integral form which not only has a structural basis but also has the property that the bilinear form on the primitive elements is self dual over Z?

2 Can the integral no-ghost theorem 6.5 be extended to show that the root spaces of non-primitive vectors are self dual, and in particular is the bilinear form on m self dual? We have shown in theorem 6.5 that it is self dual on the root spaces of roots that are either primitive or of norm 0. This question may be irrelevant, because it is possible to construct an integral form of the fake monster Lie algebra with a self dual bilinear form using the BRST complex. Over the complex numbers this construction is described in [F-G-Z] and [L-Z]. The construction can be done over the integers, and the BRST complex turns out to be a vertex superalgebra isomorphic to the vertex superalgebra of the odd self dual lattice I_{26,1}. The fake monster Lie algebra is given by the homology groups of the self adjoint BRST operator Q modulo torsion, and the homology modulo torsion of a complex with a self dual bilinear form is automatically self dual under the induced form. However it is not clear that this self dual integral form can be extended to a smooth Hopf algebra over Z.

3 Which elements of m_{α} are liftable, and in particular are all elements liftable? 4 We can also ask all the questions above about the monster Lie algebra rather than the

fake monster Lie algebra. In this case much less is known than for the fake monster Lie algebra. In particular I do not even know of an integral self dual form of the monster vertex algebra or the monster Lie algebra (but see [B-R] for some speculation about this).

5 The arguments in section 4 of [B98] proving the modular moonshine conjectures for p ≥ 13 are rather messy and computational, mainly because of the lack of a good description of what happens at the root spaces of roots α divisible by p. Is is possible to clean up this argument using the (conjectured) existence of a nice Hopf algebra for the monster Lie algebra?

References.

[A] E. Abe, Hopf algebras. Cambridge Tracts in Mathematics, 74. Cambridge University

Press, Cambridge-New York, 1980. xii+284 pp. ISBN: 0-521-22240-0

[B86] R. E. Borcherds, Vertex algebras, Kac-Moody algebras, and the monster. Proc. Natl. Acad. Sci. USA. Vol. 83 (1986) 3068–3071.

[B90] R. E. Borcherds, The monster Lie algebra, Adv. Math. Vol. 83, No. 1, Sept. 1990.

[B98] R. E. Borcherds, Modular Moonshine III, to appear in Duke Math Journal. [B-R] R. E. Borcherds, A. J. E. Ryba, Modular Moonshine II, Duke Math Journal Vol. 83

No. 2, 435-459, 1996.

[D] J. Dieudonné, Groupes de Lie et hyperalgèbres de Lie sur un corps de caractéristique p > 0 (V), Bull. Soc. Math. France 84 1956 207–239 reprinted in J. Dieudonné, “Choix d’œuvres mathématiques. Tome II.” Hermann, Paris, 1981, ISBN: 2-7056-5923-4, p. 600–632.

[D72] E. J. Ditters, Curves and formal (co)groups, Inv. Math. 17 (1972) 1-20. [F-G-Z] I. Frenkel, H. Garland, G. J. Zuckerman, Semi-infinite cohomology and string theory.

Proc. Nat. Acad. Sci. U.S.A. 83 (1986), no. 22, 8442–8446. [G-T] P. Goddard and C. B. Thorn, Compatibility of the dual Pomeron with unitarity and

the absence of ghosts in the dual resonance model, Phys. Lett., B 40, No. 2 (1972), 235-238.

[H] M. Hazewinkel, “Formal groups and applications”. Pure and Applied Mathematics, 78. Academic Press, Inc. [Harcourt Brace Jovanovich, Publishers], New York-London, 1978. ISBN: 0-12-335150-2 [K] B. Kostant, Groups over Z. 1966 Algebraic Groups and Discontinuous Subgroups

(Proc. Symposium. Pure Math., Boulder, Colorado., 1965) pp. 90–98 Amer. Math. Soc., Providence, R.I.

[L-Z] B. Lian, G. J. Zuckerman, Moonshine cohomology. Moonshine and vertex operator

algebra (Kyoto, 1994). Sūrikaisekikenkyūsho Kōkyūroku No. 904 (1995), 87–115.

[M] I. G. Macdonald, Symmetric functions and Hall polynomials. Second edition. Oxford Mathematical Monographs. Oxford Science Publications. The Clarendon Press, Oxford University Press, New York, 1995. x+475 pp. ISBN: 0-19-853489-2

[R] A. J. E. Ryba, Modular Moonshine?, In “Moonshine, the Monster, and related topics”, edited by Chongying Dong and Geoffrey Mason. Contemporary Mathematics, 193. American Mathematical Society, Providence, RI, 1996. 307-336.

[Sh] P. B. Shay, An obstruction theory for smooth formal group structure, Preprint, Hunter college, CUNY.